Grade Upto college levelGeneral Physics

The main scale reading of a screw gauge is 8 mm and the circular scale coincides with 25th division of the main scale. If the pitch of the screw gauge is 0.5 mm and the number of circular scale division is 100, then the observed reading will be

Navneet kumar
12 Years agoGrade Upto college level

1 Answer

Saurabh Singh
12 Years ago
Lc= pitch/CSD
=0.5/100
0.005mm

CSV=0.005*25

Reading=MSV+CSV
=8+.005*25

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur